OpenStudy (phebe):

MEDALSS!!! @~Jane_Doe~ 4.The area of the garden is given by the trinomial g^2-2g-24. The garden 's length is g+4. What is the garden's width? A-g-2 B-g-6 C-g - 8 D- g + 2

OpenStudy (esshotwired):

@jigglypuff314

OpenStudy (phebe):

@wolfe8

OpenStudy (anonymous):

Start by multiplying (g+4) times (g-2), if that doesn't equal g^2-2g-24, then try the next one.

OpenStudy (phebe):

ok thanx but can u giv mehh the direct answer becuz im kinda in a hurry k

OpenStudy (anonymous):

Practicing these is probably the best way to get faster at them, and thus, decrease your hurry.

OpenStudy (phebe):

ok ok well i think its C ami right

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Try looking at just the last terms. The first ones will always be g^2, so that doesn't tell you much. But you know the last term is a product of the constants.

OpenStudy (phebe):

ohh yeaaa so its A

OpenStudy (anonymous):

not quite :) (q+4)(q-2) = q^2 +2q - 8 not the g^2-2g-24 that you were looking for :) guess again?

OpenStudy (phebe):

C then

OpenStudy (phebe):

i meant D

OpenStudy (anonymous):

no, not D either (q+4)(q+2) = q^2 + 6q + 8

OpenStudy (phebe):

i think its B my cousin told mehh

OpenStudy (anonymous):

lol yes, but you have completely guessed. do you understand what FOIL is? :)

OpenStudy (phebe):

yesss FIRST OUTSIDE INSIDE LAST :) are u in connexus

OpenStudy (anonymous):

no, I'm not in connexus :P